OpenStudy (empty):

Looking for different ways to find the diagonal between the two furthest corners on a cube just for fun. :D

OpenStudy (empty):

The weirder the better.

OpenStudy (michele_laino):

 vector sum, I think

OpenStudy (michele_laino):

@Astrophysics help!

OpenStudy (michele_laino):

here is another formula, which uses the vector product and scalar product, between the sides of the cube treated as vectors: \[\Large D = \sqrt {3{l^2}} = \sqrt {3{V^{2/3}}} = \sqrt {3\left\{ {\left( {{\mathbf{l}} \times {\mathbf{l}}} \right) \cdot {\mathbf{l}}} \right\}} \] where \(V\) is the volume, and \(D\) is the diagonal, furthermore \(l\) is the side of the cube or the magnitude of the vector \(\mathbf{l}\)

OpenStudy (michele_laino):

oops.. I have made a typo: \[\Large D = \sqrt {3{l^2}} = \sqrt {3{V^{2/3}}} = \sqrt {3{{\left\{ {\left( {{\mathbf{l}} \times {\mathbf{l}}} \right) \cdot {\mathbf{l}}} \right\}}^{2/3}}} \]